OpenStudy (anonymous):

A gas of 190 mL at a pressure of 74 atm can be expected to change its pressure when its volume changes to 30.0 mL. Express its new pressure in units of atmospheres.

OpenStudy (anonymous):

if anyone could help i could sure use it

OpenStudy (abmon98):

P1*V1=P2*V2 74*190=30*P2 P2=74*190/30=468.7atm